OpenStudy (anonymous):

a boat travels 10 miles upstream in 2 hours. Going downstream it can travel 30 miles in the same time. Find the speed of the boat in still water. what would the formula look like and how would i solve formula

OpenStudy (anonymous):

let the speed of boat in still water be x let speed of stream be y

OpenStudy (anonymous):

when we go up stream the speed gets subtracted: upstream speed is: x-y

OpenStudy (anonymous):

when we go down stream the speed gets added: down stream speed is: x+y

OpenStudy (anonymous):

we know speed is distance/time use this you will get 2 equations solve them simultaneously

OpenStudy (anonymous):

so what numbers would i put where?

OpenStudy (anonymous):

30-10/x+y?

OpenStudy (anonymous):

nopes

OpenStudy (anonymous):

10/2=(x-y)

OpenStudy (anonymous):

30/2=(x+y)

OpenStudy (anonymous):

its that

OpenStudy (anonymous):

i hope you get it now

OpenStudy (anonymous):

but that would give us the answer of 10/4 no 5

OpenStudy (anonymous):

just let me check

OpenStudy (anonymous):

what answer did you get as 10/4

OpenStudy (anonymous):

speed of boat is 10

OpenStudy (anonymous):

speed of stream will be 5

OpenStudy (anonymous):

some calculation error must be there. check it

hero (hero):

d =rt and r = (b±c) 10 = (b-c)2 30 = (b+c)2 10/2 = b-c 30/2 = b+ c 5 = b-c 15 = b + c 20 = 2b 20/2 = b 10 = b 15 - 10 = c 5 = c Therefore b = 10 c = 5

OpenStudy (anonymous):

i did and both times i came out to different equations the 10/4 was a typing error i meant to say 20/4 which would give u have

OpenStudy (anonymous):

ya so thats 5 anf will give you speed of stream